# Groups with et0l co-word problem

### v1

@@ 6 Closure properties of special co-et0l-groups
The property of being a special co-et0l-group is independent of the choice of finite generating set. Furthermore, the class of special co-et0l-groups is closed under passing to finitely generated subgroups, finite direct products, and finite extensions.
@@
Let $G=\langle A\rangle$ be a special co-et0l-group and $H=\langle B\rangle\leq G$ with $B$ finite and closed under inversion. For each $b\in B$, we let $w_{b}$ be a specific word in $A^{*}$ with $b=_{G}w_{b}$. The special cspda $\mathcal{M}_{H}$ accepting $\mathsf{coWP}(H,B)$ is based on the special cspda $\mathcal{M}_{G}$ for $\mathsf{coWP}(G,A)$ but with input alphabet $B$ and with some extra states and non-reading moves. On reading a letter in $B$ in its deterministic second phase, the special cspda $\mathcal{M}_{H}$ behaves in the same way as $\mathcal{M}_{G}$ behaves on reading the word $w_{B}$. Note that by choosing $H=G$ this also proves the final assertion in the theorem.
Now let $G=H\times K$ with $H$ and $K$ special co-et0l-groups, where $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ are accepted by special cspda $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We construct a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A_{H}\cup A_{K})$ as follows. It starts with an $\epsilon$-move into the initial state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, and thereafter behaves like $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ respectively. In the first case it ignores all input letters from $A_{K}$, and in the second case it ignores those from $A_{H}$.
[…]
@@
The (standard) restricted wreath product $H\wr K$ of a special co-et0l group $H$ by a finitely generated virtually free top group $K$ is a co-et0l group.
@@
This is similar to the proof of the last part of [9, Proposition 9]. Let $H=\langle A_{H}\rangle$ be a special co-et0l-group and $\mathcal{M}_{H}$ a special cspda with $\mathcal{L}(\mathcal{M}_{H})=\mathsf{coWP}(H,A_{H})$. Let $K$ be finitely generated and virtually free. As in the proof of Theorem 3.1 above, we can choose a generating set $A_{K}=B\cup T^{\pm 1}$ of $K$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $K$ and $T$ is a right transversal of $F$ in $K$ with the element in $F$ removed, and then every element of $K$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$. As we saw in the proof of Theorem 3.1, the word problem $\mathsf{WP}(K,A_{K})$ is the language of a pushdown automaton $\mathcal{P}_{K}$ that works by storing the normal form word $wt$ of the prefix of the input word read so far, and accepts by empty stack.
@@
Let $G$ be the restricted wreath product $H\wr K$. Then $G=\langle A\rangle$ with $A=A_{H}\cup A_{K}$. (As usual, we assume that $A_{H}$ and $A_{K}$ are disjoint.) We shall describe a special cspda $\mathcal{M}_{G}$ with $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. As in [9, Proposition 9] (and earlier results on closure under restricted wreath product), it works by non-deterministically choosing a component of the base group $H^{K}$ of $G$, and using $\mathcal{M}_{H}$ to attempt to prove the non-triviality of that component of an element of the base group. A word will be accepted by $\mathcal{M}_{G}$ if and only either the input word does not project onto the identity element of $K$, or $\mathcal{M}_{H}$ accepts the word representing the chosen component of the base group.
@@
The initialisation of $\mathcal{M}_{G}$ consists of an initialisation of $\mathcal{M}_{H}$, followed by the writing of an arbitrary normal form word $wt$ on the check stack (with $t$ on top), on top of the initialised stack for $\mathcal{M}_{H}$. An arbitrary number of padding symbols p are then added to the check stack to help prevent the second extra condition of a special cspda being violated. This word $wt$ defines the component of the base group that we shall check for non-triviality using the initialisation of $\mathcal{M}_{H}$. The current state of $\mathcal{M}_{H}$ is stored at all times as a component of the state of $\mathcal{M}_{G}$.
@@
After reading the first input symbol $x$, the read write head of $\mathcal{M}_{G}$ ascends to the top of the check stack of $\mathcal{M}_{H}$ (while pushing padding symbols to the pushdown stack) and then copies the word $wt$ from the check stack to the pushdown stack. Then it passes $x$ to $\mathcal{P}_{K}$ for processing using its pushdown stack (which is the upper part of the pushdown stack of $\mathcal{M}_{G}$).
@@
Thereafter, the automaton $\mathcal{P}_{K}$ processes input symbols $x$ in this way. But whenever the pushdown stack of $\mathcal{P}_{K}$ is empty and an input symbol $x\in B$ is read, the read write head descends while removing padding symbols from the pushdown stack, and $x$ is processed by $\mathcal{M}_{H}$ using its current stored state. After $\mathcal{M}_{H}$ has processed $x$, the read write head moves back up and, unless $\mathcal{M}_{G}$ is in its entry configuration, control is passed back to $\mathcal{P}_{K}$ for processing the next input letter.
There is one further feature that is necessary for $\mathcal{M}_{G}$ to ignore the word problem of $G$. Suppose that at some reading configuration $\mathcal{M}_{H}$ is in its entry state, and the height of the pushdown stack of $\mathcal{P}_{K}$ is the same as the height of the check stack minus the padding symbols at the top of the check stack. Then $\mathcal{M}_{G}$ checks whether it should be in its entry configuration by checking whether the symbols read on the two stacks are the same and, if so, popping from the pushdown stack. It repeats this process while the symbols read are the same. If at some stage before it reaches the bottom of the pushdown stack of $\mathcal{P}_{K}$ the read symbols are not the same, then the input word does not project onto the identity element of $K$, so $\mathcal{M}_{G}$ is not in its entry configuration, and any popped symbols are replaced. If, however, this process results in the pushdown stack of $\mathcal{P}_{K}$ being emptied, then the read write head descends to the bottom of the stack of $\mathcal{M}_{G}$, and $\mathcal{M}_{G}$ returns to its entry configuration. After reading the next letter (if any), the check stack is reconstituted before processing that letter.
@@
The accept states of $\mathcal{M}_{G}$ are all reading states other than the entry states. It is now straightforward to verify that the cspda that we have described is special. Property (3) follows from the assumption that $\mathcal{M}_{H}$ has this property. ∎
@@
The class of special co-et0l-groups is closed under taking free products with finitely many free factors.
@@
It is enough to prove that the free product $G=H*K$ of two special co-et0l-groups $H$ and $K$ is a special co-et0l-group. Let $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ be special cspda accepting $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ respectively. We shall define a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A)$ with $A=A_{H}\cup A_{K}$ (which we assume to be a disjoint union). The cspda $\mathcal{M}_{G}$ includes disjoint copies of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. Beyond the single initial state of $\mathcal{M}_{G}$, each other state of $\mathcal{M}_{G}$ is naturally associated with a state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$. However, since $\mathcal{M}_{G}$ needs to do some additional processing while remembering its current state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, that current state will be stored only as the first component of a state of $\mathcal{M}_{G}$, which will also contain some further information. We denote by $\Delta_{H}$ and $E_{H}$ the stack alphabet and set of entry states for $\mathcal{M}_{H}$, and by $\Delta_{K}$ and $E_{K}$ the stack alphabet and set of entry states for $\mathcal{M}_{K}$.
@@
In the first phase (i.e. the initialisation) of $\mathcal{M}_{G}$ an arbitrary string $\alpha_{1}\beta_{1},\cdots\alpha_{k}\beta_{k}$ for some $k\geq 1$ is put onto the check stack (with $\alpha_{1}$ at the bottom), with substrings $\alpha_{i}$ and $\beta_{i}$ over $\Delta_{H}\cup E_{H}$ and $\Delta_{K}\cup E_{K}$ respectively, which are all non-empty for $i=1,\ldots,k$, except possibly for $\alpha_{1}$ and/or $\beta_{k}$ (since the string could start or end with a substring over either $\Delta_{H}\cup E_{H}$ or $\Delta_{K}\cup E_{K}$). Each non-empty string $\alpha_{i}$ or $\beta_{i}$ consists of a string over $\Delta_{H}$ or $\Delta_{K}$ that is valid for the initialisation of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ followed by the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that results from that initialisation. The check stack of $\mathcal{M}_{G}$ that we create during this initialisation is seen to be a sequence of check substacks that are initialised stacks of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We let $q_{\mathsf{ent}}$ be the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that is at the top of the lowest of those check substacks (corresponding to the first non-empty substack); we store $q_{\mathsf{ent}}$ as a component of the state of $\mathcal{M}_{G}$.
@@
After the check stack has been set, the read write head of $\mathcal{M}_{G}$ descends to the bottom of the pair of stacks. Then the first component of the entry state of $\mathcal{M}_{G}$ is set to be the entry state $q_{\mathsf{ent}}$ that was recorded at the top of the bottom check substack. This concludes the first (initialisation) phase.
@@
At all times during the second phase, exactly one of the subgroups $H$ and $K$ will be the current subgroup. When the current subgroup is $H$, the cspda $\mathcal{M}_{G}$ will have a state of $\mathcal{M}_{H}$ as its first component, and its read write head will be scanning a check substack of $\mathcal{M}_{H}$, which we call the current check stack, and similarly for $K$. The general idea is that, when the current subgroup is $H$ and input symbols from $A_{H}$ are being read, the cspda behaves in the same way as $\mathcal{M}_{H}$, using the current check substack. We call such a sequence of moves of $\mathcal{M}_{G}$ a sub-computation in $\mathcal{M}_{H}$ or in $\mathcal{M}_{K}$. As we shall see later, the cspda needs to remember the entry state of the current sub-computation, and we can store this as the second component of the current state. At the start of the second phase, the read write head is at the bottom of the stacks of $G$, the current subgroup is $H$ or $K$, depending on whether the first check substack contains $\alpha_{1}$ or $\beta_{1}$, and the current state of $\mathcal{M}_{G}$ (i.e. the entry state of this computation) is $(q_{\mathsf{ent}},q_{\mathsf{ent}})$.
@@
We need to describe how $\mathcal{M}_{G}$ behaves when it reads a letter $x$ from the input tape. Suppose first that the current subgroup is $H$. Now suppose also that $x\in A_{K}$, so $x$ is not in the current subgroup. In this case, $\mathcal{M}_{G}$ moves to the top of that current check substack while pushing padding symbols p onto the pushdown stack. At the top of this check substack, it records the current state of $\mathcal{M}_{H}$ on the pushdown stack (in the same position as the current entry state is stored on the check substack). If there are no further check substacks above the current one, then the initialisation stage of $\mathcal{M}_{G}$ was too short and $\mathcal{M}_{G}$ fails; we say that the initialisation was inadequate. Otherwise, the read write head moves up into the next check substack, which will be for the group $K$. The read write head first moves to the top of this check substack (while pushing padding symbols onto the pushdown stack) to retrieve the entry state $q_{\mathsf{ent}}$ of this initialisation of $\mathcal{M}_{K}$. Then it descends again to the bottom of this check substack (popping from the pushdown stack as it does so) and behaves as though $\mathcal{M}_{K}$ had read the input symbol $x$ while in entry state $q_{\mathsf{ent}}$. It changes to the state $(q,q_{\mathsf{ent}})$, where $q$ is the state into which a sub-computation on $\mathcal{M}_{K}$ would transition from $q_{\mathsf{ent}}$, given input $x$ and the current stack symbols, and pushes to the pushdown stack, accordingly.
@@
If, after completing a reading move and any subsequent non-reading moves of a sub-computation in the subgroup $H$, the cspda is in the entry configuration $q_{\mathsf{ent}}$ of the current sub-computation within $\mathcal{M}_{H}$ (that is, if the current state is $(q_{\mathsf{ent}},q_{\mathsf{ent}})$), but not in the entry configuration of $\mathcal{M}_{G}$ (so the read write head is at the bottom of its current check substack, but not at the bottom of the check stack of $\mathcal{M}_{G}$), then the current subgroup changes from $H$ to $K$, and the read write head descends into the check substack immediately below its current position, which will be for the group $K$. This lower check substack has recorded the configuration of a sub-computation within $\mathcal{M}_{K}$ that had been interrupted when a letter from $A_{H}$ was read. The read write head retrieves the entry state and the state of that sub-computation at the point of interruption from the tops of the check substack and pushdown stack, respectively, and then it descends and removes any padding symbols p from the top of the pushdown stack. After doing that, the configurations of both $\mathcal{M}_{G}$ and of the sub-computation in $\mathcal{M}_{K}$ are the same as they were at the point of interruption. Note that, if the next input letter is from $A_{H}$ then $\mathcal{M}_{G}$ immediately goes up again and resumes the current sub-computation in $\mathcal{M}_{H}$, but this temporary change of current subgroup and descent into the lower check substack is necessary to ensure that $\mathcal{M}_{G}$ ignores the word problem of $G$.
Reversing the roles of $H$ and $K$ in the paragraph above deals with the case where the cspda enters an entry configuration of the current sub-computation within $\mathcal{M}_{K}$ but is not in an entry configuration of $\mathcal{M}_{G}$. That completes the description of the cspda $\mathcal{M}_{G}$.
We shall show next that, so long as $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ are special then so is $\mathcal{M}_{G}$. It should be clear that $\mathcal{M}_{G}$ satisfies the additional properties (0), (1) and (2). It remains to verify property (3), that $\mathcal{M}_{G}$ ignores the word problem of $G$. So let $C$ be an entry configuration of $\mathcal{M}_{G}$ and let $uv\in A^{*}$ with $v=_{G}1$, and assume also that $C^{uv}$ does not fail. We need to show that $C^{u}=C^{uv}$.
Suppose first that the letters of $v$ are all in the same alphabet $A_{H}$ or $A_{K}$ we may assume without loss that $v\in A_{H}^{*}$. Then $v=_{H}1$, and the claim follows essentially from the assumption that $\mathcal{M}_{H}$ ignores the word problem of $H$, because $\mathcal{M}_{G}$ is carrying out a sub-computation in $\mathcal{M}_{H}$ while reading $v$. There are some complications however.
@@
It remains to prove that $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. Since the entry states of $\mathcal{G}$ are not accept states, and $\mathcal{M}_{G}$ ignores the word problem of $G$, no word $w\in A^{*}$ with $w=_{G}1$ can be accepted by $\mathcal{M}_{G}$, and so we have $\mathcal{L}(\mathcal{M}_{G})\subseteq\mathsf{coWP}(G,A)$.
@@
    $C^{w}$ is an accept state of $\mathcal{M}_{G}$; and

### v2

@@ 6 Closure properties of special co-et0l groups
The property of being a special co-et0l group is independent of the choice of finite generating set. Furthermore, the class of special co-et0l groups is closed under passing to finitely generated subgroups, finite direct products, and finite extensions.
@@
Let $G=\langle A\rangle$ be a special co-et0l group and $H=\langle B\rangle\leq G$ with $B$ finite and closed under inversion. For each $b\in B$, we let $w_{b}$ be a specific word in $A^{*}$ with $b=_{G}w_{b}$. The special cspda $\mathcal{M}_{H}$ accepting $\mathsf{coWP}(H,B)$ is based on the special cspda $\mathcal{M}_{G}$ for $\mathsf{coWP}(G,A)$ but with input alphabet $B$ and with some extra states and non-reading moves. On reading a letter in $B$ in its deterministic second phase, the special cspda $\mathcal{M}_{H}$ behaves in the same way as $\mathcal{M}_{G}$ behaves on reading the word $w_{B}$. Note that by choosing $H=G$ this also proves the final assertion in the theorem.
Now let $G=H\times K$ with $H$ and $K$ special co-et0l groups, where $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ are accepted by special cspda $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We construct a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A_{H}\cup A_{K})$ as follows. It starts with an $\epsilon$-move into the initial state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, and thereafter behaves like $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ respectively. In the first case it ignores all input letters from $A_{K}$, and in the second case it ignores those from $A_{H}$.
Now suppose that $H$ is a special co-et0l group with $\mathsf{coWP}(H,B)$ accepted by the special cspda $\mathcal{M}_{H}$, and that $H\leq G$ with $|G:H|$ finite, and let $A$ be an inverse-closed generating set of $G$ with $B\subseteq A$. Let $T$ be a right transversal of $H$ in $G$ with $1_{G}\in T$. For each $t\in T$ and $a\in A$, we have $Hta=Ht_{a}$ for some $t_{a}\in T$, and we let $w_{t,a}$ be a fixed word in $B^{*}$ with $ta=_{G}w_{t,a}t_{a}$.
@@
The (standard) restricted wreath product $H\wr K$ of a special co-et0l group $H$ by a finitely generated virtually free top group $K$ is a special co-et0l group.
@@
This is similar to the proof of the last part of [11, Proposition 9]. Let $H=\langle A_{H}\rangle$ be a special co-et0l group and $\mathcal{M}_{H}$ a special cspda with $\mathcal{L}(\mathcal{M}_{H})=\mathsf{coWP}(H,A_{H})$. Let $K$ be finitely generated and virtually free. As in the proof of Theorem 3.1 above, we can choose a generating set $A_{K}=B\cup T^{\pm 1}$ of $K$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $K$ and $T$ is a right transversal of $F$ in $K$ with the element in $F$ removed, and then every element of $K$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$. As we saw in the proof of Theorem 3.1, the word problem $\mathsf{WP}(K,A_{K})$ is the language of a pushdown automaton $\mathcal{P}_{K}$ that works by storing the normal form word $wt$ of the prefix of the input word read so far, and accepts when the stack is empty.
@@
Let $G$ be the restricted wreath product $H\wr K$. Then $G=\langle A\rangle$ with $A=A_{H}\cup A_{K}$. (As usual, we assume that $A_{H}$ and $A_{K}$ are disjoint.) We shall describe a special cspda $\mathcal{M}_{G}$ with $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. As in [11, Proposition 9] (and earlier results on closure under restricted wreath product), it works by non-deterministically choosing a component of the base group $H^{K}$ of $G$, and using $\mathcal{M}_{H}$ to attempt to prove the non-triviality of that component of an element of the base group. A word will be accepted by $\mathcal{M}_{G}$ if and only if either the input word does not project onto the identity element of $K$, or $\mathcal{M}_{H}$ accepts the word representing the chosen component of the base group.
@@
In this construction we set the bottom of stack symbol $\mathfrak{b}_{G}$ of $\mathcal{M}_{G}$ to be the bottom of stack symbol $\mathfrak{b}_{H}$ of $\mathcal{M}_{H}$. The initialisation of $\mathcal{M}_{G}$ consists of an initialisation of $\mathcal{M}_{H}$, followed by the writing of an arbitrary normal form word $wt\in A_{K}^{*}$ on the check stack (with $t$ on top), on top of the initialised stack for $\mathcal{M}_{H}$. An arbitrary number of padding symbols ${\sf p}_{G}$ followed by $\mathfrak{t}_{G}$ are then added to the check stack to help ensure that Condition (P3) is satisfied. This word $wt$ defines the component of the base group that we shall check for non-triviality using the initialisation of $\mathcal{M}_{H}$. The current state of $\mathcal{M}_{H}$ is stored at all times as a component of the state of $\mathcal{M}_{G}$.
@@
After writing $\mathfrak{b}_{G}$ at the bottom of the pushdown stack and reading the first input symbol $x$, the read write head ascends to the top of the substack of the $\mathcal{M}_{G}$ check stack that is the check stack of $\mathcal{M}_{H}$ (while pushing padding symbols ${\sf p}_{G}$ to the pushdown stack) and then copies the word $wt$ from the check stack to the pushdown stack.
At this point, and also when processing subsequent input symbols $x$ when the read write head is at the top of the pushdown stack, if $x\in A_{K}$, then $x$ is passed to $\mathcal{P}_{K}$ for processing using its pushdown stack (which is the upper part of the pushdown stack of $\mathcal{M}_{G}$). If $x\in A_{H}$ and the pushdown stack of $\mathcal{P}_{K}$ is non-empty, then $x$ is ignored by $\mathcal{M}_{G}$.
@@
If, however, the pushdown stack of $\mathcal{P}_{K}$ is empty and an input symbol $x\in A_{H}$ is read, then the read write head descends while removing padding symbols ${\sf p}_{G}$ from the pushdown stack, and $x$ is processed by $\mathcal{M}_{H}$ using its current stored state. After $\mathcal{M}_{H}$ has processed $x$, the read write head moves back up (while putting symbols ${\sf p}_{G}$ on the pushdown stack).
In any case, after processing the input letter as just described, before entering another reading state, we first check whether the cspda should be in an entry configuration, which is the case if and only if
  • (i)
    the normal form word for the element of $K$ stored at the top of the pushdown stack is equal to $wt$; and
  • (ii)
    the cspda $\mathcal{M}_{H}$ is in an entry configuration.
To see whether the first of these conditions hold, we first check whether the read write head is pointing at the top symbol $t$ of $wt$ (or to the top letter in $B$ or to $\mathfrak{t}_{H}$ if $t=\epsilon$): if not, then the two words have different lengths, and the condition does not hold. If so, then $\mathcal{M}_{G}$ checks the two words for equality by popping symbols from the pushdown stack as long as they are equal to the corresponding symbols on the check stack. If at any time the two symbols seen are not equal, then the words are unequal, and any popped symbols are replaced on the pushdown stack. If the two words are equal, then we also check whether we are in an entry configuration of $\mathcal{M}_{H}$ by checking whether it is in an entry state and, if so, checking whether the pushdown stack of $\mathcal{M}_{H}$ contains only padding symbols above $\mathfrak{b}_{H}$, by popping padding symbols ${\sf p}_{G}$. If we are not in an entry configuration of $\mathcal{M}_{H}$, then again we reconstitute all of the previous contents of the pushdown stack of $\mathcal{M}_{G}$. But if we are in an entry configuration of $\mathcal{M}_{H}$, then $\mathcal{M}_{G}$ transitions into the corresponding entry configuration of $\mathcal{M}_{G}$. This process is necessary for $\mathcal{M}_{G}$ to ignore the word problem of $G$.
At this stage we are in a reading state, either in an entry configuration of $\mathcal{M}_{G}$, or with the read write head at the top of the pushdown stack of $\mathcal{M}_{H}$, and we read the next input symbol if any. If we are in an entry configuration, then we proceed as we did for the first input letter read. We are then in any case at the the top of the pushdown stack, and we proceed as before.
@@
Following (P7), the accept states of $\mathcal{M}_{G}$ are all reading states other than the entry states. So all words that do not project onto $1_{K}$ are accepted together with all words for which the selected component in the base group has been accepted by $\mathcal{M}_{H}$ as a word in the coword problem of $H$. It is now straightforward to verify that the cspda that we have described is special. Property (P6) follows from the assumption that $\mathcal{M}_{H}$ has this property. ∎
@@
The class of special co-et0l groups is closed under taking free products with finitely many free factors.
@@
It is enough to prove that the free product G=HKG=H*K of two special co-et0l groups $H$ and $K$ is a special co-et0l group. Let $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ be special cspda accepting $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ respectively. We shall define a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A)$ with $A=A_{H}\cup A_{K}$ (which we assume to be a disjoint union). The cspda $\mathcal{M}_{G}$ includes disjoint copies of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. Beyond the single initial state of $\mathcal{M}_{G}$, each other state of $\mathcal{M}_{G}$ is naturally associated with a state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$. However, since $\mathcal{M}_{G}$ needs to do some additional processing while remembering its current state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, that current state will be stored only as the first component of a state of $\mathcal{M}_{G}$, which will have a second component containing further information. We denote by $\Delta_{H}$ and $E_{H}$ the stack alphabet and set of entry states for $\mathcal{M}_{H}$, and by $\Delta_{K}$ and $E_{K}$ the stack alphabet and set of entry states for $\mathcal{M}_{K}$.
@@
In the first phase (i.e. the initialisation) of $\mathcal{M}_{G}$ an arbitrary string $\alpha_{1}\beta_{1},\cdots\alpha_{k}\beta_{k}$ for some $k\geq 1$ is put onto the check stack (with $\alpha_{1}$ at the bottom), with substrings $\alpha_{i}$ and $\beta_{i}$ over $\Delta_{H}\cup E_{H}$ and $\Delta_{K}\cup E_{K}$ respectively, which are all non-empty for $i=1,\ldots,k$, except possibly for $\alpha_{1}$ and/or $\beta_{k}$ (since the string could start or end with a substring over either $\Delta_{H}\cup E_{H}$ or $\Delta_{K}\cup E_{K}$). With one exception, each non-empty string $\alpha_{i}$ or $\beta_{i}$ consists of a string over $\Delta_{H}$ or $\Delta_{K}$ that is valid for the initialisation of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ followed by the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that results from that initialisation. The exception is that, for the first non-empty string (which may be $\alpha_{1}$ or $\beta_{1}$) at the bottom of the check stack, we put the bottom of stack marker $\mathfrak{b}_{G}$ in place of $\mathfrak{b}_{H}$ or $\mathfrak{b}_{K}$ at the bottom. This ensures that Property (P3) is satisfied by $\mathcal{M}_{G}$.
@@
Ignoring the above exception and the symbols for the entry states of the components, we see that the check stack of $\mathcal{M}_{G}$ that we create during this initialisation is a sequence of check substacks that are initialised stacks of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We let $q_{\mathsf{ent}}$ be the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that is at the top of the lowest of those check substacks (corresponding to the first non-empty substack). After the check stack has been set, the read write head of $\mathcal{M}_{G}$ descends to the bottom of the pair of stacks, the symbol $\mathfrak{b}_{G}$ is written on the pushdown stack, and the entry state of $\mathcal{M}_{G}$ is set to $(q_{\mathsf{ent}},q_{\mathsf{ent}})$.
@@
At all times during the second phase, exactly one of the subgroups $H$ and $K$ will be the current subgroup. When the current subgroup is $H$, the cspda $\mathcal{M}_{G}$ will have a state of $\mathcal{M}_{H}$ as its first component, and its read write head will be scanning a substack that is a check stack of $\mathcal{M}_{H}$, which we call the current check stack, and similarly for $K$. The general idea is that, when the current subgroup is $H$ and input symbols from $A_{H}$ are being read, the cspda behaves in the same way as $\mathcal{M}_{H}$, using the current check substack. We call such a sequence of moves of $\mathcal{M}_{G}$ a sub-computation in $\mathcal{M}_{H}$ or in $\mathcal{M}_{K}$. As we shall see later, the cspda needs to remember the entry state of the current sub-computation, and we can store this as the second component of the current state. At the start of the second stage, the current subgroup is $H$ or $K$, depending on whether the first check substack contains $\alpha_{1}$ or $\beta_{1}$.
@@
We need to describe how $\mathcal{M}_{G}$ behaves when it reads a letter $x$ from the input tape. Suppose first that the current subgroup is $H$. Now suppose also that $x\in A_{K}$, so $x$ is not in the current subgroup. In this case, $\mathcal{M}_{G}$ moves to the top of the current check substack while pushing padding symbols ${\sf p}_{G}$ onto the pushdown stack. At the top of this check substack, it records the current state of $\mathcal{M}_{H}$ on the pushdown stack (in the same position as the current entry state is stored on the check substack). If there are no further check substacks above the current one, then the initialisation stage of $\mathcal{M}_{G}$ was too short and $\mathcal{M}_{G}$ fails; we say that the initialisation was inadequate. Otherwise, the read write head moves up into the next check substack, which will be for the group $K$. The read write head first moves to the top of this check substack (while pushing padding symbols ${\sf p}_{G}$ onto the pushdown stack) to retrieve the entry state $q_{\mathsf{ent}}$ of this initialisation of $\mathcal{M}_{K}$. Then it descends again to the bottom of this check substack (popping from the pushdown stack as it does so), pushes $\mathfrak{b}_{K}$ onto the pushdown stack, and then behaves as though $\mathcal{M}_{K}$ had read the input symbol $x$ while in entry state $q_{\mathsf{ent}}$. It changes to the state $(q,q_{\mathsf{ent}})$, where $q$ is the state into which a sub-computation on $\mathcal{M}_{K}$ would transition from $q_{\mathsf{ent}}$, given input $x$ and the current stack symbols, and pushes to the pushdown stack, accordingly.
@@
Suppose that, after completing a reading move and any subsequent non-reading moves of a sub-computation in the subgroup $H$, the cspda is in the entry configuration of the current sub-computation within $\mathcal{M}_{H}$ but not in the entry configuration of $\mathcal{M}_{G}$. Note that this situation arises if and only if the current state is $(q_{\mathsf{ent}},q_{\mathsf{ent}})$, where $q_{\mathsf{ent}}$ is the entry state of the current computation, and the current symbol seen on the check stack is the bottom of stack marker $\mathfrak{b}_{H}$ (and not $\mathfrak{b}_{G}$). Then the current subgroup changes from $H$ to $K$, and the read write head descends into the check substack immediately below its current position, which will be for the group $K$. This lower check substack has recorded the configuration of a sub-computation within $\mathcal{M}_{K}$ that had been interrupted when a letter from $A_{H}$ was read. The read write head retrieves the entry state and the state of that sub-computation at the point of interruption from the tops of the check substack and pushdown stack, respectively, and then it descends and removes any padding symbols ${\sf p}_{G}$ from the top of the pushdown stack. After doing that, the configurations of both $\mathcal{M}_{G}$ and of the sub-computation in $\mathcal{M}_{K}$ are the same as they were at the point of interruption. Note that, if the next input letter is from $A_{H}$ then $\mathcal{M}_{G}$ immediately goes up again and resumes the current sub-computation in $\mathcal{M}_{H}$, but this temporary change of current subgroup and descent into the lower check substack is necessary to ensure that $\mathcal{M}_{G}$ ignores the word problem of $G$.
Reversing the roles of $H$ and $K$ in the paragraph above deals with the case where the cspda enters an entry configuration of the current sub-computation within $\mathcal{M}_{K}$ but is not in an entry configuration of $\mathcal{M}_{G}$. Following (P7), the accepting states of $\mathcal{M}_{G}$ are all reading states other than the entry states. That completes the description of the cspda $\mathcal{M}_{G}$.
We shall show next that, so long as $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ are special then so is $\mathcal{M}_{G}$. It should be clear that $\mathcal{M}_{G}$ satisfies the additional properties (P1)–(P5) and (P7). It remains to verify Property (P6), that $\mathcal{M}_{G}$ ignores the word problem of $G$. So let $C$ be an entry configuration of $\mathcal{M}_{G}$ and let $uv\in A^{*}$ with $v=_{G}1$, and assume also that $C^{uv}$ does not fail. We need to show that $C^{u}=C^{uv}$.
Suppose first that the letters of $v$ are all in the same alphabet $A_{H}$ or $A_{K}$ - we may assume without loss that $v\in A_{H}^{*}$. Then $v=_{H}1$, and the claim follows essentially from the assumption that $\mathcal{M}_{H}$ ignores the word problem of $H$, because $\mathcal{M}_{G}$ is carrying out a sub-computation in $\mathcal{M}_{H}$ while reading $v$. There are some complications however.
@@
It remains to prove that $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. Since the entry states of $\mathcal{M}_{G}$ are not accepting states, and $\mathcal{M}_{G}$ ignores the word problem of $G$, no word $w\in A^{*}$ with $w=_{G}1$ can be accepted by $\mathcal{M}_{G}$, and so we have $\mathcal{L}(\mathcal{M}_{G})\subseteq\mathsf{coWP}(G,A)$.
@@
    $C^{w}$ is an accept configuration of $\mathcal{M}_{G}$; and